OpenStudy (anonymous):

Alex spent 3/7 of his money. He gave 1/4 of the remainder to his sister. He had $120 left. How much money did he have in the beginning?

OpenStudy (phi):

you should change the English into "math" start with How much money did he have in the beginning? call it x (unknown amount that he started with) he spent 3/7 of his money. replace "his money" with x: 3/7 of x which is written as 3/7 * x he started with x, and spent 3/7 x : \( x - \frac{3}{7} x \) is how much he has left gave 1/4 of the remainder to his sister. 1/4 of the remainder is \( \frac{1}{4}\left(x - \frac{3}{7} x\right) \) \[ x - \frac{3}{7} x -\frac{1}{4}\left(x - \frac{3}{7} x\right) = 120\]

OpenStudy (anonymous):

Oh! I get it now. Thanks a bunch! :]